Circular No. 6271

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995bb IN ANONYMOUS GALAXY S. Tokarz and P. Garnavich, Center for Astrophysics, report their dis-

covery of a supernova from the spectrum of an anonymous galaxy observed for the 15R Redshift Survey directed by M. Geller: “The galaxy (magnitude R ∼ 15.0 via M. Kurtz) is an edge-on spiral or irregular located at α = 0^{h}16^{m}17^{s}.64, δ = +12^{o}24′53′′.4 (equinox 2000.0). A spectrum of the galaxy, taken by P. Berlind on Nov. 29 using the Mt. Hopkins 1.5-m Tillinghast telescope, shows a mixture of narrow and broad emission lines. The unresolved lines are identified as H II emission at z = 0.0058. The most intense broad line is measured at 634.7 nm and is identified as [O I]. Other, weaker emission features appear at 458.7, 593.7, and 734.8 nm and correspond to Mg I], Na I, and [Ca II], respectively. The overall spectrum is similar to that of the type-Ib supernova 1985F, ∼ 300 days after maximum (Schlegel and Kirshner 1989, A.J. 98, 577). The FWHM of the [O I] emission is 6600 km/s, while the Mg I] line has a FWHM of 4600 km/s. The visual absorption is 0.2–0.3 mag, based on the narrow-line Balmer decrement. Assuming a Hubble parameter of 65 km s^{−1} Mpc^{−1}, the [O I] 630.0-nm/636.3-nm luminosity is estimated to be 6 × 10^{39} erg/s. This implies a large oxygen mass, comparable to that of SN 1985F. CCD images obtained by C. Prosser with the Mt. Hopkins 1.2-m telescope on Dec. 7 shows the supernova coincident with the weak galaxy nucleus. The supernova may have reached m_{pg} ∼ 13 early in 1995, and prediscovery observations are needed to determine the time of maximum.”

OJ 287 M. Kidger, Instituto de Astrofisica de Canarias, reports that OJ 287

is passing through an extended double-peaked outburst in very close agreement with the date predicted by Lehto and Valtonen (Ap.J., in press) for a binary black-hole model. Recent data show that this source is almost as bright as during the maximum a year ago: 1995 Nov. 10.2 UT, V = 14.59 (P. Boltwood, Ottawa, Canada); 25.080, 14.73 (Y. Efimov, Crimea, Ukraine); Dec. 2.40, 15.09 (Boltwood); 3.92, 14.8 (L. Takalo, Tuorla, Finland); 4.25, 14.61 (Kidger); 5.27, 14.22 (T. Balonek, Colgate Obs., communicated by J. Webb); 8.25, 14.28 (Takalo and K. Nilsson); 8.26, 14.27 (Boltwood). M. Valtonen comments that the lightcurve’s “double-peak structure has been obvious in previous superflares as well, but in 1983 and 1984 the first peak was more prominent, while in 1994 the second peak was higher.”

1995 December 8 Daniel W. E. Green